Costs of peanuts is 2.00 per ound and the cost of raisins is 3.50 per pound, how many pounds of peanuts and raisins shoul be used to make a 3 pound mixture that costs 3.00 per pound
Answer by checkley77(12844)

3.5x+2(3-x)=3*3
3.5x+6-2x=9
1.5x=9-6
1.5x=3
x=3/1.5
x=2 pounds od raisons are used.
3-2=1 pound of peanuts is used.
Proof:
3.5*2+2*1=9
7+2=9
9=9